# DecoupledESC: Enhancing Emotional Support Generation via Strategy-Response Decoupled Preference Optimization

### v1

@@ 4.2 Datasets Statistics
Model Type Flu.$\uparrow$ Pro.$\uparrow$ Emp.$\uparrow$ Hel.$\uparrow$
Qwen Chosen 3.82 3.52 3.20 2.95
Rejected 3.65 3.09 2.41 2.32
Improve ($\uparrow$) 4.66% 13.92% 32.78% 27.16%
Llama Chosen 3.99 3.74 3.33 3.09
Rejected 3.93 3.21 2.40 2.39
Improve ($\uparrow$) 1.53% 16.52% 38.75% 29.29%
@@ 4.3 Datasets Quality
We evaluate the content quality of $100$ samples from $D_{\text{RG-dpo}}$ using gpt-4.1-mini-2025-04-14222https://openai.com/index/gpt-4-1. As shown in Table 2, chosen responses outperform rejected ones across four LLM-based metrics, with over a $30\%$ gain in Empathy.
Backbone Paradigm Method Automatic Metrics.$\uparrow$ LLM-based Metrics.$\uparrow$ Strategy Metrics.
@@
Table 3: Comparison of models under different optimization paradigms and training methods. The best score is in-bold, while the second best score is underlined.
@@ 5.1 Decoupled ESC Framework
To address the issue raised in Obs 13.1) and Obs 23.2), that vanilla training of strategy selection and response generation can lead to negative optimization, hindering the reduction of preference bias and the improvement of response quality. As shown in Figure 5, we propose a decoupled ESC optimization framework, inspired by the Extended Process Model of Emotion Regulation (EPMER) Gross (2015), which divides emotion regulation into three sequential stages: identification, strategy selection, and implementation. We decouple the ESC generation process into two independent subtasks: Strategy Planning and Response Generation. This enables more stable and controllable training for each.
@@ 5.2 Decoupled-SFT and Decoupled-DPO
To further optimize the Strategy Planner and Response Generator and reduce psychological errors, we apply the offline reinforcement learning method (DPO). Leveraging the previously constructed preference datasets $D_{\text{SP-dpo}}$ and $D_{\text{RG-dpo}}$, we separately train the two modules to improve their decision-making and response generation capabilities. For the Strategy Planner, we train the model on the strategy preference dataset $D_{\text{SP-dpo}}$ to enhance its preference for gold strategies while mitigating bias toward suboptimal strategies, thereby reducing strategy selection bias. The loss function is defined as follows:
@@
For the Response Generator, we train with $D_{\text{RG-dpo}}$ to enhance the quality and empathy of responses through the following loss function:
$\mathcal{L}_{\text{RG-dpo}}=-\mathbb{E}_{(c,s,a_{c},a_{r})\sim D_{\text{RG-dpo}}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(a_{c}|c,s)}{\pi_{\text{ref}}(a_{c}|c,s)}\right.\right.$
Through this decoupled two-stage DPO optimization, we effectively reduce bias in strategy planning and enhance the empathy of responses, resulting in more robust and emotionally intelligent dialogue generation.
@@ 6 Experiments
  • RQ1: Can Vanilla-DPO better mitigate preference bias and improve response quality compared to Vanilla-SFT?
  • RQ2: Does the decoupled ESC framework lead to better bias reduction and response quality?
  • RQ3: In the decoupled setting, is DPO more effective than SFT?
  • RQ4: Can Decoupled-DPO more effectively reduce common psychological errors by separately optimizing strategy and response?
@@ 6.1 Experimental Setup
Backbone Models.
We conducted experiments using two widely adopted large language models: Qwen2.5-7B-Instruct Team (2024) and Llama3.1-8B-Instruct Dubey et al. (2024).
@@
We use a learning rate of 1e-5 and a batch size of 32 for SFT, training for 3 epochs, while DPO is trained for 1 epoch. Experiments are conducted on 4×24GB RTX4090 GPUs. For LLM-based evaluation, we randomly sample 100 instances from the test set and assess them using gpt-4.1-mini-2025-04-1422footnotemark: 2. More implementation details are provided in Appendix D.1.
@@ 6.2 Experimental Results
To answer RQ1, whether DPO can mitigate preference bias and improve response quality in vanilla setting, we compare Vanilla-DPO with Vanilla-SFT. As shown in Table 3, DPO consistently outperforms SFT on LLM-based metrics for both Qwen and Llama, indicating enhanced response quality.
However, preference evaluation results are mixed: for Qwen, DPO slightly reduces the preference bias ($\mathcal{B}$) compared to SFT; for Llama, DPO shows higher bias and lower prediction accuracy. We attribute this to DPO’s greater sensitivity to data quality in the vanilla setting. When training data contains conflicting optimization signals between strategy and content, DPO may suffer from negative transfer, leading to performance degradation. To further examine how different preference data affect coupled models, we provide a detailed comparison in Appendix B.
@@
Decoupled vs. Vanilla (RQ2).
To answer RQ2, whether the decoupled ESC framework better mitigates preference bias and improves empathetic response quality. We compare Decoupled and Vanilla paradigms under both SFT and DPO training. As shown in Table 3, decoupled models consistently outperform their vanilla counterparts across most metrics, including automatic metrics, LLM-based metrics, and strategy metrics. Notably, the preference bias of Decoupled-DPO models trained with Qwen and Llama are $0.22$ and $0.15$, respectively, representing a substantial improvement over Vanilla-DPO and Vanilla-SFT. These results highlight the effectiveness of decoupled optimization in mitigating preference bias. Furthermore, we invited 20 master’s students with a background in psychology to evaluate 20 samples under the guidance of a licensed psychological counselor. As shown in Figure 6, Decoupled-DPO outperforms Vanilla-DPO across all four LLM-based metrics.
We attribute this to the decoupled framework, which avoids the conflict and negative optimization often encountered in vanilla training. It also simplifies the construction of preference datasets. In addition, the decoupled strategy planning module can serve as an external planner, which, as Kang et al. Kang et al. (2024) note, helps reduce strategy bias and improves the coherence and empathy of responses.
@@
To answer RQ3, whether DPO outperforms standard SFT in a decoupled setting, we compare Decoupled-DPO and Decoupled-SFT. As shown in Table 3, Decoupled-DPO outperforms other models on most LLM-based and bias metrics for both Qwen and Llama, achieving state-of-the-art (SOTA) performance. Importantly, the improvements of DPO over SFT are more pronounced in the decoupled setting, with bias for Qwen dropping from $0.27$ to $0.22$, and for Llama from $0.23$ to $0.15$. In contrast, the vanilla framework shows little change from SFT to DPO, further validating the synergy between decoupling and DPO training. Table 4 illustrates the impact of DPO training on the Response Generator. Given the ground truth strategy, the DPO-trained Response Generator consistently outperforms the SFT model across all LLM-based metrics.
Backbone GT_Stra SFT DPO Flu.$\uparrow$ Pro.$\uparrow$ Emp.$\uparrow$ Hel.$\uparrow$
@@
Table 4: Ablation study on Response Generator.
@@
To verify the effectiveness of Decoupled-DPO in improving response quality, we adopt the same error categorization approach as Obs 1 (§3.1). As shown in Figure 2 and 7, Decoupled-DPO achieves the highest proportion of No Error cases at 27%, outperforming Qwen-SFT and Llama-SFT by an average of 7%. This demonstrates that Decoupled-DPO effectively reduces common psychological errors and improves overall response quality.
@@ 7 Conclusion
In this paper, we propose a Decoupled ESC framework that separates strategy planning from empathetic response generation, enabling targeted optimization and avoiding mutual interference. Extensive experiments demonstrate that our Decoupled ESC framework significantly outperforms joint optimization baselines, effectively reducing preference bias and improving response quality in ESC tasks.
@@
Due to computational resource constraints, our study validates the proposed method only on relatively small-scale models (7B, 8B, and 9B). While significant improvements are observed at this scale, the effectiveness on larger models (e.g., 70B) remains an open question and is a key direction for future work. In addition, the generalizability of our decoupled ESC optimization framework warrants further investigation under other preference optimization methods such as KTO, SimPO, and IPO.
@@
This research utilizes the ESConv and FailedESConv dataset Liu et al. (2021), which has been obtained with proper authorization and in compliance with data usage agreements. We ensure that all data used in this study are handled responsibly and in accordance with ethical standards, respecting the privacy and confidentiality of individuals involved. All necessary agreements and permissions for the use of this dataset have been signed, ensuring full compliance with data protection regulations.
@@
It should be noted that while the model demonstrates certain capabilities in psychological support tasks, the strategies and expressions it adopts differ from those used in real-life professional counseling. Given the diversity of users’ emotional states and circumstances, the model’s responses may not always align with professional therapeutic standards and, in some cases, may unintentionally affect users’ emotional well-being. Therefore, this model is intended for academic research purposes only and is not recommended for commercial use. Caution is advised when applying the model outside controlled research environments, and it should not be used for real-world psychological counseling without supervision by qualified professionals.
@@ A.1 Definitions of Gross’s Extended Process Model of Emotion Regulation
The Extended Process Model of Emotion Regulation, proposed by Gross in 2015 Gross (2015), refines earlier models by conceptualizing emotion regulation as a temporally ordered process comprising three core stages: Identification, Selection, and Implementation.
@@ A.2 Definitions of Psychological Errors
Under the guidance of a licensed psychological counselor and based on psychological literature Raskin and Rogers (2005); Stebnicki (2007), we identified common errors frequently made by counselors in real-world therapy sessions. These were categorized into five types of empathy-related psychological errors:
@@ A.3 Definitions of Counseling Stages
Backbone Chosen Rejected Automatic Metrics.$\uparrow$ LLM-based Metrics.$\uparrow$ Strategy Metrics.
@@ E.2 LLM Metrics Criteria
Table 8 summarizes the LLM evaluation metrics, including Fluency, Professionalism, Empathy, and Helpfulness, along with their descriptions, evaluation criteria, and scoring scales. All metrics are rated on a 5-point Likert scale Joshi et al. (2015). Specifically, Fluency and Empathy are adapted from the ESC-Eval framework Zhao et al. (2024), Professionalism is guided by the CPsyCoun framework Zhang et al. (2024a), and Helpfulness is derived from the SoulChat evaluation setup Chen et al. (2023).
@@ E.3 Prompt for LLM Metrics
Figure 171819, and 20 present the prompts used for LLM-based evaluation of Fluency, Professionalism, Empathy, and Helpfulness, respectively. Each prompt explicitly defines the role of the LLM as a judge and outlines the corresponding evaluation criteria. To minimize potential bias, the prompts are carefully designed to avoid revealing model names or being influenced by text length.
@@ E.4 Human Evaluation
To complement the LLM-based evaluation and enhance the credibility of our results, we conducted a human evaluation involving 20 master’s students with a background in psychology, under the guidance of a licensed psychological counselor, on 20 samples generated by our Decoupled-DPO model based on the Llama backbone. As shown in Figure 6, the Decoupled-DPO achieves superior performance compared to Vanilla-DPO across multiple evaluation metrics.

### v2

@@ 4.3 Datasets Quality
We evaluate the content quality of $100$ samples from $D_{\text{RG-dpo}}$ using gpt-4.1-mini-2025-04-14444https://openai.com/index/gpt-4-1. As shown in Table 3, chosen responses outperform rejected ones across four LLM-based metrics, with over a $30\%$ gain in Empathy.
[…]
@@
Table 4: Comparison of models under different optimization paradigms and training methods. The best score is in-bold, while the second best score is underlined. $\uparrow$ means a higher score is better whereas $\downarrow$ is exactly the opposite.
@@ 5.1 Decoupled ESC Framework
To address the issue raised in Obs 1 and Obs 2, that vanilla training of strategy planning and response generation can lead to negative optimization, hindering the reduction of preference bias and the improvement of response quality. As shown in Figure 5, we propose a Decoupled ESC optimization framework, inspired by the Extended Process Model of Emotion Regulation (EPMER) Gross (2015), which divides emotion regulation into three sequential stages: Identification, Strategy Selection, and Implementation. We decouple the ESC generation process into two independent subtasks: Strategy Planning and Response Generation. This enables more stable and controllable training for each.
@@ 5.2 Decoupled-SFT and Decoupled-DPO
To reduce psychological errors, we further optimize the Strategy Planner and Response Generator using the offline RL method (DPO). Based on the preference dataset IPM-PrefDial, which includes $D_{\text{SP-dpo}}$ and $D_{\text{RG-dpo}}$, we separately train both modules to enhance strategy selection and response generation. For the Strategy Planner, we apply DPO on $D_{\text{SP-dpo}}$ to encourage preference for gold strategies and reduce bias toward suboptimal ones. The loss function is defined as:
@@
For the Response Generator, we apply DPO on $D_{\text{RG-dpo}}$ to improve response quality and empathy, with the loss function defined as:
[…]
@@ 6 Experiments
  • RQ1: What are the performance differences between DPO and SFT in bias mitigation and response generation within a coupled framework?
  • RQ2: What specific advantages does decoupling strategy planning and response generation bring to ESC tasks?
  • RQ3: In the Decoupled ESC framework, how do SFT and DPO respectively affect the model’s bias and generation quality?
  • RQ4: To what extent can the Decoupled ESC framework effectively reduce psychological errors?
@@ 6.1 Experimental Setup
Backbones.
We conducted experiments using two LLMs: Qwen2.5-7B-Instruct Team (2024) and Llama3.1-8B-Instruct Dubey et al. (2024).
@@
For SFT, we train for 3 epochs using a batch size of 32 and a learning rate of 1e-5. DPO is trained for 1 epoch under the same settings. All experiments are conducted on 4×24GB RTX 4090 GPUs. For LLM-based evaluation, we randomly sample 100 test instances and evaluate them using gpt-4.1-mini-2025-04-1444footnotemark: 4 alongside 3 psychology experts. Additional implementation details are provided in Appendix D.1.
[…]
Vanilla-SFT 15.75 18.11 3.31 2.52 2.22 2.06 0.26
Decoupled-SFT 16.97 19.12 3.87 2.74 2.39 1.95 0.23
Vanilla-DPO 15.15 17.25 3.41 2.79 2.41 2.28 0.28
Decoupled-DPO 17.50 19.16 3.81 2.73 2.64 2.17 0.15
Table 5: Results of SFT and DPO under Vanilla and Decoupled Paradigms for Llama3.1-8B-Instruct.
@@ 6.2 Experimental Results
As shown in Table 4, under the vanilla setting, DPO consistently outperforms SFT on LLM-based metrics for both Qwen and Llama, indicating enhanced response quality.
However, Strategy-Metrics evaluation shows mixed results: DPO slightly reduces bias ($\mathcal{B}$) on Qwen but increases bias and lowers accuracy on Llama. We attribute this to DPO’s sensitivity to noisy data in the vanilla setting—conflicting optimization signals between strategy and content can cause negative transfer (see section Obs 2). We further compare the impact of different preference data on coupled models in Appendix B.
Decoupled vs. Vanilla (RQ2).
As shown in Tables 4 and 5, decoupled models outperform vanilla ones on most metrics. Notably, the preference bias of Decoupled-DPO models drops to $0.22$ (Qwen) and $0.15$ (Llama), much better than Vanilla-DPO and Vanilla-SFT. This shows decoupled optimization effectively reduces preference bias.
Human Evaluation
[…]
@@
We attribute this to the decoupled framework. Its strategy planning module, as shown by Kang et al. Kang et al. (2024), acts as an external planner that helps reduce strategy preference bias. It also avoids the optimization conflict in vanilla training and simplifies preference data construction.
@@
As shown in Tables 4 and 5, DPO brings greater improvements over SFT in the decoupled setting, with bias reduced from $0.27$ to $0.22$ for Qwen, and from $0.23$ to $0.15$ for Llama. In contrast, the vanilla framework yields minimal bias reduction from SFT to DPO, highlighting the stronger synergy between decoupling and DPO training. Table 6 further shows that, given ground-truth strategies, the DPO-trained Response Generator consistently outperforms its SFT counterpart across all LLM-based metrics.
[…]
@@
Table 6: Ablation study on Response Generator. ✓ indicates that the method or data is used in training, while ✗ indicates it is not. GT: Ground Truth Strategy.
@@
To verify the effectiveness of Decoupled-DPO in improving response quality, we adopt the same error categorization approach as Obs 1. By contrasting Figures 2 and 7, we observe that Decoupled-DPO achieves the highest proportion of No Error cases at 2727%, outperforming Qwen-SFT and Llama-SFT by an average of 77%. This demonstrates that Decoupled-DPO effectively reduces common psychological errors and improves overall response quality.
@@ 7 Conclusion
In this paper, we propose a Decoupled ESC framework that separates strategy planning from empathetic response generation, enabling targeted optimization and avoiding mutual interference. Extensive experiments demonstrate that our Decoupled ESC framework significantly outperforms joint optimization baselines, effectively reducing preference bias and improving response quality in Emotional Support Conversation tasks.
@@
Our study, while demonstrating promising results, has several limitations that suggest avenues for future research. The annotation of psychological errors relied on an in-context learning approach to manage the resource-intensive nature of expert labeling; while efficient, this method’s classifications may not be perfectly consistent with expert judgment. Furthermore, our experiments were constrained by computational resources to models up to 9B parameters, so the efficacy of our method on much larger models (e.g., 70B) requires future validation. The generalizability of our decoupled framework also warrants broader investigation, as it was primarily tested with DPO and should be assessed with other algorithms like KTO, SimPO, and IPO. Finally, this work is confined to textual analysis, and a key future direction is to extend our framework to incorporate multimodal inputs such as speech and facial expressions for more holistic emotional support.
@@
This research utilizes the ESConv and FailedESConv dataset Liu et al. (2021), which has been obtained with proper authorization and in compliance with data usage agreements. We ensure that all data used in this study is handled responsibly and in accordance with ethical standards, respecting the privacy and confidentiality of individuals involved. All necessary agreements and permissions for the use of this dataset have been signed, ensuring full compliance with data protection regulations.
@@
It should be noted that while the model demonstrates certain capabilities in psychological support tasks, its strategies cannot encompass the full range of approaches and techniques used in real-life professional counseling. Given the diversity of users’ emotional states and circumstances, the model’s responses may not always align with professional standards and may unintentionally affect users’ emotional well-being. Therefore, this model is intended for academic research only and is not recommended for commercial use. Caution is advised when using it beyond research settings, and it should not be applied to real-world counseling without professional supervision.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (62293554, U2336212), “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2024C01073), Ningbo Innovation “Yongjiang 2035” Key Research and Development Programme (2024Z292), and Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
@@ A.1 Definitions of Gross’s Extended Process Model of Emotion Regulation
The Extended Process Model of Emotion Regulation (EPMER), proposed by Gross in 2015 Gross (2015), refines earlier models by conceptualizing emotion regulation as a temporally ordered process comprising 3 core stages:
@@ A.2 Definitions of Psychological Errors
Under the guidance of 3 psychological experts and based on psychological literature Raskin and Rogers (2005); Stebnicki (2007), we identified common errors frequently made by psychological experts in real-world therapy sessions. These were categorized into 5 types of empathy-related psychological errors [^5] All definitions of psychological errors were reviewed by 3 psychological experts.:
@@ A.3 Definitions of Counseling Stages
[…]
@@ E.2 LLM Metrics Criteria
Table 10 summarizes the LLM evaluation metrics, including Fluency, Professionalism, Empathy, and Helpfulness, along with their descriptions, evaluation criteria, and scoring scales. All metrics are rated on a 5-point Likert scale Joshi et al. (2015). Specifically, Fluency and Empathy are adapted from the ESC-Eval framework Zhao et al. (2024), Professionalism is guided by the CPsyCoun framework Zhang et al. (2024), and Helpfulness is derived from the SoulChat evaluation setup Chen et al. (2023).
@@ E.3 Prompt for LLM Metrics
Figures 171819, and 20 present the prompts used for LLM-based evaluation of Fluency, Professionalism, Empathy, and Helpfulness, respectively. Each prompt explicitly defines the role of the LLM as a judge and outlines the corresponding evaluation criteria. To minimize potential bias, the prompts are carefully designed to avoid revealing model names or being influenced by text length.
@@ E.4 Human Evaluation
To complement the LLM-based evaluation and enhance the credibility of our results, we conducted a human evaluation with 3 licensed psychology experts on 100 samples generated by our Decoupled-DPO model based on the Llama backbone. In addition, we performed a pairwise win/loss comparison between Decoupled-DPO and Vanilla-DPO (both using Llama) on another set of 100 samples, with inter-rater agreement again measured by Fleiss’ Kappa ($\kappa$) Fleiss and Cohen (1973). As shown in Figure 6, Decoupled-DPO consistently outperforms Vanilla-DPO across all four LLM-based metrics. The evaluator agreement, measured by $\kappa$, further supports the reliability of the human judgments.